For Binomial distribution with n = 10 and probability of success p = 0.4
find probability of less than 3 successes (x < 3) (3 is not included).
Use Appendix Table for Binomial distribution.

Answer by ewatrrr(24785)

P(x < 3) = 1-P(x ≥ 3) = 1 - (1-P(x ≤2)) = .1673